----------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------------1.)given the following equation y''+4y=f(x) find the general solution y(x) in terms of the continuous function f(x) , which is not a sine or cosine function
Answer #1

Therefore, the general solution is given by You = c.ft P.y You = ce the in sein in fronde 14 en fain fantan ain -in + I 41 42page2y+4y= foxes -0 Auxiliary equation of given differential equation is m²+4=0 (but y= m² 3 mati bet mi=21, M2=-2i Thus, the copage1